OpenStudy (bekkahmaclean):

a catapult launches a boulder with an upward velocity of 122 ft/s. the height of the boulder, (h), in feet after t seconds is given by the function h(t)--16t^2+122t+10. how long does it take the boulder to reach its maximum height? whats the boulder's maximum height? round to the nearest hundredth, if necessary.

OpenStudy (bekkahmaclean):

reaches a maximum height of 15.42 feet after 7.71 seconds reaches a maximum height of 7.71 after 3.81 seconds reaches a maximum height of 242.56 feet after 7.62 seconds reaches a maximum height of 242.56 feet after 3.81 seconds

OpenStudy (welshfella):

one way would be to convert the formula to vertex form by completing the square

OpenStudy (welshfella):

-16t^2+122t+10 =-16( t^2 - 7.625t) + 10 = -16[ {t - 3.8125)^2 - 3.8125^2)] + 10 = -16(t - 3.8125) ^2 + 232.5625 + 10 = -16(t - 3.81)^2 + 242.56

OpenStudy (welshfella):

- so the time is 3.81 and maximum height is 242.56

OpenStudy (welshfella):

vertex form of a quadratic is a(x - b^2 + c where the vertex is at (b , c)

OpenStudy (bekkahmaclean):

oh man thank you so mcuh i was struggling so hard with that one